# On the Relationship between Space-Time Accessibility and Leisure Activity Participation

### v1

@@ Abstract
Understanding how accessibility shapes participation in leisure activities is central to promoting inclusive and vibrant urban life. Conventional accessibility measures often focus on potential access from fixed home locations, overlooking the constraints and opportunities embedded in daily routines. In this study, we introduce a space–time accessibility (SPA) metric rooted in the capability approach, capturing feasible leisure opportunities between home and work given a certain time budget, individual transport modes, and urban infrastructure. Using high-resolution GPS data from 2,415 residents in the Paris region, we assess how SPA influences total travel time and leisure participation, measured as the diversity of leisure activity locations. Spatial patterns show that most individuals—especially active transport users—choose destinations aligned with their SPA-defined opportunity sets, underscoring the metric’s validity in capturing capability sets. Structural equation modeling reveals that SPA directly fosters leisure diversity but also reduces travel time, which in turn is associated with lower diversity. These findings highlight the value of person-centered, capability-informed accessibility metrics for understanding inequalities in urban mobility and informing transport planning strategies that expand real freedoms to participate in social life across diverse population groups.
@@ 1 Introduction
Transport accessibility is central to inclusive cities through its influence on individual mobility and activity participation (Allen and Farber, 2020; Luz et al., 2022; Liao et al., 2025a). By shaping how easily people can reach different locations and activities, spatial accessibility defines the scope of places and social environments individuals can access (Pereira et al., 2017), particularly to those third-place activities that foster meaningful social interactions (Oldenburg and Brissett, 1982). When transport opportunities are limited, the risk of social exclusion rises for certain groups (Luz and Portugal, 2022; Gallego Méndez et al., 2023). Studies have shown that greater proximity to certain destinations is associated with higher participation rates in those activities (e.g., Althoff et al., 2025). Yet, it remains unclear how transport accessibility, when accounting for individual constraints, links to participation in third-place activities and how this varies across population groups.
Visits to third places shape individuals’ social exposure by linking them to wider society beyond institutional settings, with the extent of such exposure often conditioned by people’s flexibility in allocating time to discretionary activities (Oldenburg and Brissett, 1982). Focusing on leisure activities, as an important part of third-place activities, draws attention to how transport systems must accommodate not only traditional trip purposes, like fixed home-work commuting trips, but also the diverse and time-sensitive mobility demands associated with rich and diverse activities that improve individuals’ well-being (Lee, 2022).
Travel times and time budgets play a central role in shaping individuals’ ability to engage in out-of-home daily activities. A longer travel time enables individuals to extend beyond proximate locations and diversify their activity choices, albeit at the cost of having less available time to engage in activities, whereas tighter time budgets restrict them to a narrower set of opportunities, often closer to home or work anchors. Activity locations in proximity reduce the required travel effort for reaching comparable opportunities, thereby lowering the effective time budget needed to satisfy activity demand. Structural constraints—the layout of the urban amenities and transportation infrastructure—can limit people’s ability to access resources outside the home, hindering efficient time use and participation in broad third-place activities (Luz and Portugal, 2022).
Such structural constraints from the built environment can be largely captured with spatial accessibility measures through network-based models of potential access (Levinson and King, 2020). Nonetheless, most studies in the literature mobilize such transport accessibility measures using home-based and static measures (Ryan et al., 2023), which largely overlook trip chaining and trips that start from non-home destinations, particularly work. Many leisure activities, central to urban vibrancy and the fostering of social interactions (Botta and Gutiérrez-Roig, 2021), take place after work rather than from home (McGuckin and Murakami, 1999). Unlike home- or work-based measures, space-time accessibility metrics account for the spatio-temporal constraints of an individual’s time–space prism (Kim and Kwan, 2003; Victoriano et al., 2020; Saraiva and Barros, 2022), offering a more realistic proxy for their ability to access leisure opportunities.
@@
To date, our understanding of how transport accessibility shapes leisure activity participation remains largely static. Constrained space–time accessibility can reduce mobility efficiency and temporal flexibility, and thereby contribute to limited leisure activity participation. In this study, we i) quantify individual space–time accessibility to leisure opportunities, ii) examine its explanatory power for total travel time and leisure activity participation (location diversity), and iii) assess how these relationships vary across population groups.
@@ 1.1 Related work
Third-place activities—such as gym, café, and park—significantly shape social interactions, identity, and community connectedness (Oldenburg and Brissett, 1982). For example, some third places, such as shopping centres, strategically located to connect diverse neighbourhoods, could attract visitors across socioeconomic groups and foster inter-group exposure (Nilforoshan et al., 2023). Institutional locations, i.e., homes and workplaces, serve as anchors that structure daily mobility patterns, constraining the accessibility and relevance of third places (Miller, 1991). The effect of these anchors on mobility is intertwined with individual attributes, lifestyle, and activity demand (Luz and Portugal, 2022).
The proposition that accessibility should be understood as a human capability (Pereira et al., 2017; Luz and Portugal, 2022) offers a unifying perspective that links the multiple factors shaping one’s potential access and observed mobility patterns. This framing consists of two interacting components, the spatial environment and resources, and the individual’s conversion factors. Spatial environment and resources include land-use patterns, transport systems, and their temporal constraints (Liu et al., 2023), while the conversion factors reflect individuals’ perceptions and abilities to translate these resources into real access, i.e., actual travel behavior and activity participation. The conversion factor is shaped by their travel time budget (Chung and Nam, 2024) and cognitive limitations, as well as the wider social, economic, and political environment (Ryan et al., 2019). The interaction of these elements can either expand or restrict people’s capabilities; when restrictions become severe (accessibility poverty), systemic barriers to opportunities lead to transport-related social exclusion (Lucas, 2012), with individuals experiencing different forms of exclusion depending on how spatial resources, personal conversion factors, and external environments combine.
@@
A growing body of evidence shows a strong link between accessibility and realized mobility. Higher accessibility levels are strongly correlated with participation in total, mandatory, and discretionary activities (Fransen et al., 2018; Luz et al., 2022). Better job accessibility is associated with greater mobility ranges (Liao et al., 2025b), and proximity to walkable areas promotes higher levels of physical activity (Althoff et al., 2025). Improving accessibility in neighborhoods with concentrations of low-income or carless households located outside major transit corridors has also been shown to increase daily activity participation (Allen and Farber, 2020). Similarly, enhancing public transit accessibility can particularly benefit discretionary activity participation among disadvantaged groups (Zhang et al., 2022).
@@
Nevertheless, most existing studies rely on home-based cumulative opportunity (Allen and Farber, 2020; Luz et al., 2022) or simple proximity measures (McCormack and Shiell, 2011) of accessibility and treat individual attributes separately. This approach is essentially static and overlooks the mechanisms through which accessibility shapes mobility. For example, while greater job accessibility may be associated with higher leisure participation at the aggregate level (Luz et al., 2022), such a misalignment between the activity types used to measure accessibility and those used to measure participation obscures policy relevance and undermines individual-level insights.
@@
In contrast, space–time accessibility is an individual-based approach that enables the examination of intra-group and intra-location heterogeneities, thereby offering a deeper understanding of accessibility inequalities (Neutens et al., 2010; Saraiva and Barros, 2022). It builds on two core concepts: an individual’s trajectory in space and time (space-time path), and the set of potential trajectories an individual could take given temporal and spatial constraints (anchors) (Kwan, 1998). They define the geographical extent of accessible opportunities (space-time prisms). Measures of space–time accessibility range from direct depictions of prisms (Miller, 1991) to empirically scalable methods based on travel time budgets (Saraiva and Barros, 2022).
Transport accessibility directly shapes travel behavior, with consequences for mobility efficiency—an intermediate factor influencing activity participation. A key expression of efficiency is trip chaining, where individuals combine multiple activities into a single journey rather than returning home between stops (Zhu and Guo, 2022). Recent advances in space–time accessibility incorporate this behavior implicitly by accounting for the feasibility of chaining activities between home and work, given an individual’s travel time budget, residential and workplace locations, and the built environment (Saraiva and Barros, 2022). This formulation reflects realistic mobility patterns, such as engaging in activities on the way home from work, and thus provides a useful proxy for assessing both mobility efficiency and opportunities for third-place participation.
@@
Structural Equation Modeling (SEM) has been widely used in travel behavior research, demonstrating its utility in analyzing complex relationships between the built environment and mobility (Golob, 2003; Kroesen and Van Wee, 2022; Song et al., 2016). SEM is particularly adept at dissecting intricate causal structures (Golob, 2003; Kroesen and Van Wee, 2022). For instance, SEM has been applied to explore the interdependencies among land use, socio-demographic attributes, and travel patterns (Song et al., 2016), and to model mediating variables such as car ownership when analyzing the built environment’s influence on travel behavior (Zhang et al., 2025). This method provides valuable insights into how built environment characteristics—including density, diversity, connectivity, and accessibility—affect various outcomes, such as individual transport emissions, health outcomes, and rural income, by shaping residents’ travel choices and capabilities (Azmoodeh et al., 2023; Kroesen and Van Wee, 2022; Li et al., 2024; Song et al., 2016).
@@ 2 Conceptual framework
In this study, we conceptualize space–time accessibility (SPA) as a human capability (Luz and Portugal, 2022), where individuals’ time budgets and mobility resources interact to shape the opportunity sets that on can reach, including third-place activities such as leisure (Figure 1). Conversion factors as a reinforcing cycle: the ability to convert spatial and temporal resources into actual activity participation (functionings) contributes to well-being, which in turn strengthens individuals’ capacity to utilize resources and expand future capabilities.
For example, a person with greater flexibility to integrate social or leisure stops into daily travel (trip chaining) can increase their well-being through richer social interaction, which may enhance confidence and mobility skills that improve their ability to access other opportunities. Conversely, limited time budgets can reduce third-place participation, narrowing social exposure and reinforcing inequalities. Transport planning approaches that equate accessibility solely with the provision of infrastructure or services often overlook these individual-level conversion processes, which are shaped by personal characteristics, environmental barriers, and cultural norms (Pereira et al., 2017). This capability-oriented framework thus shifts attention to how varying levels of space–time accessibility are necessary to ensure equal opportunities, social inclusion, and the freedom to engage in third activities essential for well-being and development.
In this study, we hypothesize that the factors shaping an individual’s participation in leisure activities are those indicated by the thick green arrows (Figure 1). The list of all opportunities accessible by a person (their capability set) is conceptualized as arising from two building blocks: individuals’ home and work/study locations, and their transport and amenity resources.
Figure 1: Accessibility as a human capability. Adapted from (Luz and Portugal, 2022; Luz et al., 2022).
@@ 3.1 Study area: The Paris region
The Paris region, commonly known as the Île-de-France, is the most populous of France’s eighteen regions, with an estimated 12.27 million residents as of January 2023 (INSEE, 2022). Centered on the capital Paris in the north-central part of the country, it covers 12,012 km2, about 2% of metropolitan France, yet accounts for nearly 20% of the national population. In the Paris region, population density reaches about 20,200 inhabitants/km2 in the city of Paris, decreases to around 7,000 inhabitants/km2 in the inner suburbs, and averages roughly 3,700 inhabitants/km2 across the wider conurbation (La Grande Conversation, 2023).
The Paris region (see Figure 2), with its monocentric urbanization pattern, has one of the world’s densest and most multifaceted transport networks—spanning multiple transit modes and innovative services (metro, RER, tramways, buses, and on-demand services), coordinated by Île‑de‑France Mobilités (IDFM). The region also offers robust ecosystem support via Mobility-as-a-Service (MaaS) platforms, integrating public transit with shared mobility, MaaS apps, and seamless ticketing (Île-de-France Mobilités, 2023). Residents feature frequent multimodal trips and rich variation across individuals (Yin and Leurent, 2023).
Figure 2: The Paris region. Black dots in the bottom-right corner of the figure indicate anchor locations within the highlighted box.
@@ 3.2 Trip records
We use a dataset collected via the Enquête Mobilité par GPS (EMG 2023) initiative and made available through participating in the NetMob 2025 Data Challenge (Chasse et al., 2025). The EMG 2023 survey, conducted between October 2022 and May 2023 covered 3,337 residents aged 16–80 in the Paris region, excluded non-residents, tourists, and the immobile, and used a multichannel recruitment strategy combining quotas and random draws to ensure representativeness (Chasse et al., 2025).
In this study, we analyze 14,169 individual-day records from 2,415 individuals with identified home and work locations, including information on date, day type (e.g., strike, holiday), origin, destination, transport mode, trip duration, and purpose. Origins and destinations are represented by the centroids of the visited locations’ corresponding H3 hexagons at resolution 10 (Uber Technologies, Inc., 2018), each covering approximately 0.015 km2.

### v2

@@ Abstract
Understanding how accessibility shapes participation in leisure activities is central to promoting inclusive and vibrant urban life. Conventional accessibility measures often focus on potential access from fixed home locations, overlooking the constraints and opportunities embedded in daily routines. In this study, we apply a space–time accessibility (STA) metric rooted in the capability approach, capturing feasible leisure opportunities between home and work given a certain time budget, individual transport modes, and urban infrastructure. Using high-resolution GPS data from 2,415 residents in the Paris region, we assess how STA influences total travel time and leisure participation, measured as the diversity of leisure locations visited. Our analysis shows that most individuals—especially active transport users—choose destinations aligned with their STA-defined opportunity sets, underscoring the metric’s validity in capturing capability sets. Structural equation modeling reveals that STA directly fosters leisure diversity but also reduces travel time, which in turn is associated with lower diversity of visited leisure locations. These findings highlight the value of person-centered, capability-informed accessibility metrics for understanding inequalities in urban mobility and informing transport planning strategies that expand real freedoms to participate in social life across diverse population groups.
@@ 1 Introduction
Transport accessibility is central to inclusive cities through its influence on individual mobility and activity participation (allen2020planning; luz2022does; liao2025socio). By shaping how easily people can reach different locations and activities, spatial accessibility defines the scope of places and social environments individuals can access (pereira2017distributive), particularly activities taking place in third places, that is, informal settings outside home and work (e.g., cafés, community centres, parks) that foster meaningful social interactions (oldenburgThirdPlace1982). When transport opportunities are limited, the risk of social exclusion rises for certain groups (luz2022understanding; gallego2023social). Studies have shown that greater proximity to certain destinations is associated with higher participation rates in those activities (e.g., reimers2014proximity). Yet, it remains unclear whether access to transportation effectively is associated with more participation in third-place activities, and how individual capabilities and time constraints shape this relationship.
Although third-place activities are crucial for social inclusion and well-being, they have received less scholarly attention than commuting and essential travel domains in accessibility literature (luz2022does; tomasiello2023unfolding). Visits to third places affect individuals’ social exposure by linking them to wider society beyond institutional settings (oldenburgThirdPlace1982). Such exposure often depends on people’s flexibility in allocating time to discretionary activities—that is, non-obligatory activities such as socializing, leisure, or informal errands that they choose to engage in. Recent studies reflect the growing attention to third-place activities in accessibility research, e.g., evaluating inequalities in access to destinations like parks (tomasiello2023unfolding; barboza2024comparative). Among the third-place activities, this paper focuses specifically on leisure activities. This choice highlights that transport systems must accommodate the diverse and time-sensitive mobility demands associated with activities that improve individuals’ well-being (lee2022third).
Accessibility to leisure opportunities is shaped by transport systems, which play a key role in enabling activity participation, yet it is usually examined in a static manner (luz2022understanding). Home- or work-based accessibility measures, typically operationalized through network-based models of potential access (levinson2020transport), estimate how many leisure destinations can be reached within a given travel-time threshold. Most studies in the literature utilize such home-based and static measures of transport accessibility (ryan2023accessibility). There are two major limitations of static measures. First, this stream largely overlooks trip chaining and trips that originate from non-home locations, particularly workplaces. Many leisure activities, central to urban vibrancy and the fostering of social interactions (botta2021modelling), take place after work rather than from home (mcguckin1999examining; farber2013social; barboza2024comparative). Second, static approaches neglect the dynamics of individuals being in motion throughout their daily activities. One of the key factors shaping individuals’ ability to engage in leisure activities is their travel time budget, defined as the amount of time people are willing to spend traveling each day. Individuals who pursue a wider range of activities—particularly discretionary ones, such as leisure, social engagements, or personal errands—tend to accumulate longer daily travel time budgets, as these additional and geographically dispersed activities extend the time spent traveling (joly2016intensive). Conversely, individuals with tighter activity schedules or fewer discretionary commitments exhibit more limited daily travel-time budgets and remain anchored to a narrower set of proximate destinations. Activity locations in proximity reduce the required travel effort for reaching comparable opportunities, thereby lowering the effective time budget needed to satisfy activity demand (chen2021effects).
Unlike home- or work-based measures, space-time accessibility metrics account for the spatio-temporal constraints of an individual’s time–space prism (kim2003space; victoriano2020time; saraiva2022accessibility), offering a more realistic proxy for their ability to access leisure opportunities. However, space–time accessibility metrics are data-intensive and vary substantially across individuals, requiring detailed information to compute, including home and work locations (saraiva2022accessibility). The resulting measures also require careful interpretation, as understanding them often depends on additional personal attributes to disentangle the complex factors shaping accessibility outcomes (chen2021effects). Consequently, applying these metrics—and studying their downstream effects on activity participation—generally requires high-resolution datasets, such as the one used in the present study.
@@
To date, our understanding of how transport accessibility shapes leisure activity participation remains largely static. This study draws on a unique geolocation dataset capturing multi-day movement trajectories for over 2,000 residents of the Greater Paris region, alongside detailed socio-demographic and household information. In this study, we i) quantify individual space–time accessibility to leisure opportunities, ii) examine its explanatory power for total travel time and leisure activity participation (location diversity), and iii) assess how these relationships vary across population groups.
The remainder of this manuscript is organized as follows. The introduction continues with a brief review of related work (Section 1.1), followed by an explanation of the applied conceptual framework (Section 2). We then describe the materials (Section 3) and methodology (Section 4), followed by Section 5, which presents the results and accompanying discussion. The paper concludes with a summary of the main findings in the Conclusion section.
@@ 1.1 Related work
Third-place activities play a central role in urban life, yet their use and relevance are strongly structured by individuals’ home and work places and personal circumstances. Third-place activities—such as gym, café, and park—significantly shape social interactions, identity, and community connectedness (oldenburgThirdPlace1982; farber2013social). For example, some third places, such as shopping centres, strategically located to connect diverse neighbourhoods, could attract visitors across socioeconomic groups and foster inter-group exposure (nilforoshan2023human). Institutional locations, i.e., homes and workplaces, serve as anchors that structure daily mobility patterns, constraining the accessibility and relevance of third places (miller1991modelling). The effect of these anchors on mobility is intertwined with individual attributes, lifestyle, and activity demand (luz2022understanding; barboza2024comparative).
@@
A growing body of evidence indicates a strong correlation between transportation accessibility and physical activity participation. Higher accessibility levels are strongly correlated with participation in total, mandatory, and discretionary activities (fransen2018spatio; luz2022does). Proximity to walkable areas promotes higher levels of physical activity (althoff2025countrywide). Improving accessibility in neighborhoods with concentrations of low-income or carless households located outside major transit corridors has also been shown to increase daily activity participation (allen2020planning). Similarly, enhancing public transit accessibility can particularly benefit participation in discretionary activities among disadvantaged groups (zhang2022eliminating).
@@
Nevertheless, most existing studies rely on home-based cumulative opportunity (allen2020planning; luz2022does) or simple proximity measures mccormack2011search of accessibility, treating individual attributes separately. This approach is essentially static and overlooks the mechanisms through which accessibility shapes mobility. For example, while greater job accessibility may be associated with higher leisure participation at the aggregate level (luz2022does), such a misalignment between the activity types used to measure accessibility and those used to measure participation obscures policy relevance and undermines individual-level insights.
@@
In contrast, space–time accessibility is a person-based approach that enables the examination of intra-group and intra-location heterogeneities, thereby offering a deeper understanding of accessibility inequalities across individuals (neutens2010equity; saraiva2022accessibility). It builds on two core concepts: an individual’s trajectory in space and time (space-time path), and the set of potential trajectories an individual could take given temporal and spatial constraints (anchors) (kwan1998space). They define the geographical extent of accessible opportunities (space-time prisms). Measures of space–time accessibility range from direct depictions of prisms (miller1991modelling) to empirically scalable methods based on travel time budgets (saraiva2022accessibility). Recent advances in space–time accessibility account for the feasibility of chaining activities between home and work, given an individual’s travel time budget, residential and workplace locations, and the built environment (saraiva2022accessibility; barboza2024comparative). This formulation reflects realistic mobility patterns, such as engaging in leisure activities on the way home from work, and thus provides a useful proxy for assessing both mobility efficiency and opportunities for third-place participation.
@@
To disentangle the complex pathways through which space–time accessibility shapes leisure participation, we employ Structural Equation Modeling (SEM), a widely used approach in travel behavior research (Golob_2003; Kroesen_Van_Wee_2022; Song_et_al_2016). SEM is particularly adept at dissecting intricate causal structures (Golob_2003; Kroesen_Van_Wee_2022). For instance, SEM has been applied to explore the interdependencies among land use, socio-demographic attributes, and travel patterns (Song_et_al_2016), and to model mediating variables such as car ownership when analyzing the built environment’s influence on travel behavior (Zhang_et_al_2025). This method has been shown to provide valuable insights into how built environment characteristics—including density, diversity, connectivity, and accessibility—affect various outcomes, such as individual transport emissions, health outcomes, and rural income, by shaping residents’ travel choices and capabilities (Azmoodeh_et_al_2023; Kroesen_Van_Wee_2022; Li_et_al_2024; Song_et_al_2016).
@@ 2 Conceptual framework
In this study, we operationalize the notion of accessibility as a human capability using the space–time accessibility (STA) framework (pereira2017distributive; luz2022understanding). In this approach, individuals’ time budgets and mobility resources (e.g., main transport mode, amenity distribution) interact to shape the opportunity sets that one can reach, including third-place activities such as leisure (Figure 1). Within this framework, the ability to convert spatial and temporal resources into actual participation (individual conversion function), particularly in social or third-place activities, plays a crucial role in overall well-being. Such participation can also contribute to improved capability of utilizing these spatial and temporal resources, even if we do not explicitly model these reinforcing effects.
Transport planning approaches that equate accessibility solely with the provision of infrastructure or services often overlook these individual-level conversion processes, which are shaped by personal characteristics, environmental barriers, and cultural norms (pereira2017distributive; van2022disentangling; luz2022understanding). This capability-oriented framework thus shifts attention to how varying levels of space–time accessibility are necessary to ensure equal opportunities, social inclusion, and the freedom to engage in third activities essential for well-being and development.
To make these concepts more intuitive, we can think of a person’s capability set as corresponding to the full range of activities they could realistically engage in, such as visiting a gym, meeting friends at a café, or visiting a park. Their individual conversion function reflects the personal and contextual factors that determine how easily they can turn accessible opportunities into actual participation. This includes socio-demographic characteristics (e.g., age, gender, income), household responsibilities (e.g., childcare obligations), mobility skills, and constraints such as fatigue or safety concerns. Finally, the resources underpinning accessibility include one’s transport options (e.g., car access, bike availability, public transit service), time budgets, and the spatial distribution of nearby amenities. Together, these elements shape not only what opportunities are theoretically reachable but also which ones a person can meaningfully engage in during daily life.
In this study, we hypothesize that the factors shaping an individual’s participation in leisure activities are the individual conversion function, the capability set, and the mobility, indicated by the thick green arrows in Figure 1. The list of all opportunities accessible by a person (their capability set) is operationalized from two building blocks: individuals’ home and work/study locations, and their transport and amenity resources.
Figure 1: Accessibility as a human capability. Conceptual framework illustrating how we conceptualize space–time accessibility (STA) as a human capability, adapted from luz2022understanding; luz2022does. The diagram shows how factors influencing activity participation interact with capability sets.
@@ 3 Materials
This study draws on a combination of large-scale mobility traces, contextual geographic data, and socio-economic information to characterize individuals’ opportunities and participation in activities in the Greater Paris region (Section 3.1). We integrate high-resolution GPS travel diaries from the 2023 Enquête Mobilité par GPS (EMG) (Section 3.2) with detailed attributes on household and personal characteristics (Section 3.3), public transport and road network infrastructure (Section 3.4), and point-of-interest (POI) distributions (Section 3.5). Together, these materials provide a comprehensive foundation for quantifying space–time accessibility and examining the determinants of leisure and third-place activity participation across diverse population groups.
@@ 3.1 Study area: The Paris region
The Paris region, commonly known as the Île-de-France, is the most populous of France’s eighteen regions, with an estimated 12.27 million residents as of January 2023 (insee_pop_2020_regions). Centered on the capital Paris in the north-central part of the country, it covers 12,012 km2, about 2% of metropolitan France, yet accounts for nearly 20% of the national population. In the Paris region, population density reaches about 20,200 inhabitants/km2 in the city of Paris, decreases to around 7,000 inhabitants/km2 in the inner suburbs, and averages roughly 3,700 inhabitants/km2 across the wider conurbation (lagrandeconversation2023).
The Paris region (see Figure 2), with its monocentric urbanization pattern, has one of the world’s densest and most multifaceted transport networks—spanning multiple transit modes and innovative services (metro, RER, tramways, buses, and on-demand services), coordinated by Île‑de‑France Mobilités (IDFM). The region also offers robust ecosystem support via Mobility-as-a-Service (MaaS) platforms, integrating public transit with shared mobility, MaaS apps, and seamless ticketing (idfmmob_mobility_as_a_service_2023). Residents frequently take multimodal trips and exhibit rich variation in their mobility patterns across individuals (yin2023multimodal).
Figure 2: The Paris region and its public transit lines. Grey lines represent IRIS boundaries, and black lines depict roads and motorways. The right panel provides a zoomed-in view of central Paris, where green points mark home and work locations.
@@ 3.2 Trip records
We use a dataset collected via the Enquête Mobilité par GPS (EMG 2023) initiative and made available through participating in the NetMob 2025 Data Challenge (netmob25). The EMG 2023 survey, conducted between October 2022 and May 2023 covered 3,337 residents aged 16–80 in the Paris region, excluded non-residents, tourists, and the immobile, and used a multichannel recruitment strategy combining quotas and random draws to ensure representativeness (netmob25).
In this study, we analyze 14,169 individual-day records from 2,415 individuals with identified home and work locations, including information on date, day type (e.g., strike, holiday), origin, destination, transport mode, trip duration, and purpose. Origins and destinations are represented by the centroids of the visited locations’ corresponding H3 hexagons at resolution 10 (uberh3), each covering approximately 0.015 km2.